# Error Broadcast and Decorrelation as a Potential Artificial and Natural Learning Mechanism

### v2

@@ Appendix I Supplementary on numerical experiments
The models were trained on an NVIDIA Tesla V100 GPU, using the hyperparameters detailed in the sections below. Each experiment was conducted five times under identical settings, and the reported results reflect the average performance. We used the standard train/test splits for the datasets, with MNIST comprising 60,000 training examples and CIFAR-10 comprising 50,000, while both datasets included 10,000 test examples. The MNIST dataset [52] is made available under the Creative Commons Attribution-Share Alike 3.0 license. The CIFAR-10 dataset [45], originating from the University of Toronto, is publicly available for academic research purposes. Both datasets were accessed via standard deep learning library functionalities. Rather than utilizing automatic differentiation tools, we manually implemented the gradient calculations for the EBD algorithm, utilizing batched operations to ensure computational efficiency. As a side note, the $(1-\lambda)$ factors present in the derived update expressions are absorbed into the learning rate constants and thus eliminated. In our experiments, we trained the MLP models for 120 epochs and the CNN and LC models for 100 epochs on MNIST and 200 epochs on CIFAR-10. In addition, we trained the CorInfoMax-EBD (3-layer, batch size = 20) model for 60 epochs.
@@ I.1 Architectures
The architectural details of MLP, CNN and LC networks for the MNIST and CIFAR-10 datasets are shown in Tables 3 and 4, respectively. The structure of the models are the same as in the reference [9]. In all architectures, we used ReLU as the nonlinear functions except the last layer. Furthermore, the architectural details of the biologically more realistic CorInfoMax network for MNIST and CIFAR-10 datasets are shown in Table 5. These techniques are the same as examples in Appendix J.5 of [41].
Table 3: MNIST architectures. FC: fully connected layer. Conv: convolutional layer. LC: locally connected layer. For fully connected layers, layer_size is shown which corresponds to the size of the hidden layer. For convolutional and locally connected layers, (num_channels, kernel_size, stride, padding) are shown.
@@
I.2 Accuracy and loss curves
Figures 5h and 6 present the training/test accuracy and MSE loss curves over epochs for the CIFAR-10 and MNIST datasets. Solid lines represent test curves; dashed lines denote training curves.
Figure 5: Train and test accuracies plotted as a function of algorithm epochs for training with various update rules (averaged over $n=5$ runs associated with the corresponding $\pm$ std envelopes) for the (a) MLP on MNIST dataset (b) MLP on CIFAR-10 dataset (c) CNN on MNIST dataset (d) CNN on CIFAR-10 dataset (e) LC on MNIST dataset (f) LC on CIFAR-10 dataset (g) 10-Layer CorInfoMax-EBD with batchsize=1 on MNIST dataset (h) 10-Layer CorInfoMax-EBD with batchsize=1 on CIFAR-10 dataset
Figure 6: Train and test mean squared errors (MSE) plotted as a function of algorithm epochs for training with various update rules (averaged over $n=5$ runs associated with the corresponding $\pm$ std envelopes) for the (a) MLP on MNIST dataset (b) MLP on CIFAR-10 dataset (c) CNN on MNIST dataset (d) CNN on CIFAR-10 dataset (e) LC on MNIST dataset (f) LC on CIFAR-10 dataset (g) 3-Layer CorInfoMax-EBD with batchsize=20 on MNIST dataset (h) 3-Layer CorInfoMax-EBD with batchsize=20 on CIFAR-10 dataset.
I.3 CorInfoMax-EBD
In this section, we offer additional details regarding the numerical experiments conducted with the CorInfoMax Error Broadcast and Decorrelation (CorInfoMax-EBD) algorithm. Appendix I.3.1 elaborates on the general implementation details. Appendix I.3.2 presents the fundamental learning steps of the algorithm, which are based on the EBD method. Appendices I.3.3 and I.3.4 discuss the initialization of the algorithm’s variables and describe the hyperparameters. Finally, Appendix I.3.5 ( 3-Layer and batch size=20), Appendix I.3.6 (3-Layer batch size=1) and Appendix I.3.7 (10-Layer batch size=1) detail the specific hyperparameter configurations used in our numerical experiments for the MNIST and CIFAR-10 datasets. In Appendix I.2 we present the accuracy and loss learning curves for the CorInfoMax-EBD, shown in Figures 5h.(g)-(h) and Figures 6.(g)-(h), respectively.
I.3.1 Implementation details
We implemented the CorInfoMax-EBD algorithm based on the repository available at:
https://github.com/BariscanBozkurt/Supervised-CorInfoMax
This repository from Bozkurt et al. [41], used as a basis for our CorInfoMax-EBD implementation, did not specify an explicit license in its public repository at the time of access. Our use and modification are for academic research purposes, building upon the published scientific work presented in [41].
This repository was referenced in [41]. The following modifications were made to the original code:
@@ I.3.2 Algorithm
1:Learning rate parameters $\lambda_{d}$,$\lambda_{E}$, $\mu^{(d,k)}[m]$, $\mu^{(f,k)}[m]$, $\mu^{(b,k)}[m]$
@@
13:${\mathbf{Q}}^{(k)}[m]\leftarrow\hat{{\mathbf{R}}}_{{\mathbf{g}}^{(k)}({\mathbf{h}}^{(k)})\bm{\epsilon}}[m]\,{\mathbf{E}}^{(k)}[m]$
@@
19:$\Delta{\mathbf{W}}^{(f,k)}_{\text{EBD}}[m]\leftarrow-\dfrac{\mu^{(d_{f},k)}[m]}{B}\,\bm{\Phi}^{(k)}[m]{{\mathbf{H}}^{(k-1)}[m]}^{\top}$
20:$\Delta{\mathbf{W}}^{(f,k)}_{\text{Pred}}[m]\leftarrow\dfrac{\mu^{(f,k)}[m]}{B}\,\overset{\rightarrow}{{\mathbf{E}}}^{(k)}[m]\,\left({\mathbf{H}}^{(k-1)}[m]\right)^{\top}$
21:${\mathbf{W}}^{(f,k)}[m]\leftarrow{\mathbf{W}}^{(f,k)}[m-1]+\Delta{\mathbf{W}}^{(f,k)}_{\text{EBD}}[m]+\Delta{\mathbf{W}}^{(f,k)}_{\text{Pred}}[m]$
@@
24:$\Delta{\mathbf{W}}^{(b,k)}_{\text{EBD}}[m]\leftarrow-\dfrac{\mu^{(d_{b},k)}[m]}{B}\,\bm{\Phi}^{(k)}[m]\,{{\mathbf{H}}^{(k+1)}[m]}^{\top}$
25:$\Delta{\mathbf{W}}^{(b,k)}_{\text{Pred}}[m]\leftarrow\dfrac{\mu^{(b,k)}[m]}{B}\,\overset{\leftarrow}{{\mathbf{E}}}^{(k)}[m]\,{\mathbf{H}}^{(k+1)}[m]^{\top}$
26:${\mathbf{W}}^{(b,k)}[m]\leftarrow{\mathbf{W}}^{(b,k)}[m-1]+\Delta{\mathbf{W}}^{(b,k)}_{\text{EBD}}[m]+\Delta{\mathbf{W}}^{(b,k)}_{\text{Pred}}[m]$
@@
29:$\Delta{\mathbf{B}}^{(k)}_{\text{EBD}}[m]\leftarrow-\dfrac{\mu^{(d_{l},k)}[m]}{B}\bm{\Phi}^{(k)}[m]\,{\mathbf{H}}^{(k)}[m]^{\top}$
30:$\Delta{\mathbf{B}}^{(k)}_{\text{E}}[m]\leftarrow-\dfrac{\gamma_{E}}{B}{\mathbf{Z}}^{(k)}[m]{\mathbf{Z}}^{(k)}[m]^{T}$
31:${\mathbf{B}}^{(k)}[m]\leftarrow\frac{1}{\lambda_{E}}{\mathbf{B}}^{(k)}[m]+\Delta{\mathbf{B}}^{(k)}_{\text{E}}[m]+\Delta{\mathbf{B}}^{(k)}_{\text{EBD}}[m]$
I.3.3 Initialization of algorithm variables
We initialize the variables ${\mathbf{W}}^{(f,k)}$, ${\mathbf{W}}^{(b,k)}$, and ${\mathbf{R}}_{{\mathbf{h}}^{(k)}\bm{\epsilon}}$ using PyTorch’s Xavier uniform initialization with its default parameters for the MNIST dataset. For the CIFAR-10 dataset is initialized with gain $0.25$. For the lateral weights ${\mathbf{B}}^{(k)}$, we first generate a random matrix ${\mathbf{J}}^{(k)}$ of the same dimensions, also using the Xavier uniform distribution, with gain$=1$ for the MNIST dataset and with gain$=0.5$ for the CIFAR-10 dataset. We then compute ${\mathbf{B}}^{(k)}[0]={\mathbf{J}}^{(k)}{{\mathbf{J}}^{(k)}}^{T}$, ensuring that ${\mathbf{B}}^{(k)}[0]$ is a positive definite symmetric matrix.
I.3.4 Description of hyperparameters
@@
I.3.5 Hyperparameters for 3-layer MNIST and CIFAR-10 datasets for batch size 20
Table 7 and 8 list the hyperparameters used in the 3-Layer CorInfoMax-EBD numerical experiments for the MNIST and CIFAR-10 datasets with a batch size of $20$, where $m$ denotes the iteration index in both.
Table 7: 3-Layer CorInfoMax-EBD hyperparameters: MNIST dataset, $20$-batch size.
Hyperparameter Value
$\alpha[m]$ $\frac{1}{3\times 10^{-3}\times\lfloor\frac{m}{10}\rfloor+1}$
$\alpha_{2}[m]$ $\frac{1}{3\times\lfloor\frac{m}{10}\rfloor+1}$
$\mu^{(d_{f},k)}[m]$ $\left[\begin{array}[]{ccc}96&60&1e5\end{array}\right]\alpha[m]$
$\mu^{(d_{b},k)}[m]$ $\left[\begin{array}[]{ccc}96&60&1e5\end{array}\right]\alpha[m]$
$\mu^{(d_{l},k)}[m]$ $\left[\begin{array}[]{ccc}0.25&0.25&0.25\end{array}\right]\alpha[m]$ for epoch$=0$
$\left[\begin{array}[]{ccc}0.5&0.5&0.5\end{array}\right]\alpha[m]$ for epoch$>0$
$\mu^{(f,k)}[m]$ $\left[\begin{array}[]{ccc}0.11e-18&0.06e-18&0.035e-18\end{array}\right]\alpha[m]$
$\mu^{(b,k)}[m]$ $\left[\begin{array}[]{cc}1.125e-18&0.375e-18\end{array}\right]\alpha[m]$
$\mu^{(p,k)}[m]$ $\left[\begin{array}[]{ccc}4.4e-3&6e-3&3.5e-12\end{array}\right]\alpha_{2}[m]$
$p^{(k)}$ $\left[\begin{array}[]{ccc}2.5&2.5&0.1\end{array}\right]$
$\mu^{(k)}_{f,\ell_{1}}[m]$ $\left[\begin{array}[]{ccc}0.008&0.135&0\end{array}\right]\alpha_{2}[m]$
$\mu^{(k)}_{b,\ell_{1}}[m]$ $\left[\begin{array}[]{ccc}0&0.35&0.05\end{array}\right]\alpha_{2}[m]$
$\mu^{(k)}_{f,w-\ell_{2}}[m]$ $\frac{8e-2}{10^{-2}\times\lfloor\frac{m}{10}\rfloor+1}$
$\mu^{(k)}_{b,w-\ell_{2}}[m]$ $\frac{8e-2}{10^{-2}\times\lfloor\frac{m}{10}\rfloor+1}$
$\lambda_{E}$ $0.999999$
$\lambda_{d}$ $0.99999$
$B$ $20$
Table 8: 3-Layer CorInfoMax-EBD hyperparameters: CIFAR-10 dataset, $20$-batch size.
Hyperparameter Value
$\alpha[m]$ $\frac{1}{3\times 10^{-3}\times\lfloor\frac{m}{10}\rfloor+1}$
$\alpha_{2}[m]$ $\frac{1}{3\times\lfloor\frac{m}{10}\rfloor+1}$
$\mu^{(d_{f},k)}[m]$ $\left[\begin{array}[]{ccc}80&50&1e5\end{array}\right]\alpha[m]$ for epoch$=0$
$\left[\begin{array}[]{ccc}320&400&1e5\end{array}\right]\alpha[m]$ for epoch$>0$.
$\mu^{(d_{b},k)}[m]$ $\left[\begin{array}[]{ccc}0&0&0\end{array}\right]\alpha[m]$
$\mu^{(d_{l},k)}[m]$ $\left[\begin{array}[]{ccc}0.5&0.5&0.5\end{array}\right]\alpha[m]$ for epoch$=0$
$\left[\begin{array}[]{ccc}2.0&2.0&2.0\end{array}\right]\alpha[m]$ for epoch$>0$
$\mu^{(f,k)}[m]$ $\left[\begin{array}[]{ccc}0.11e-18&0.06e-18&0.035e-18\end{array}\right]\alpha[m]$
$\mu^{(b,k)}[m]$ $\left[\begin{array}[]{cc}1.125e-18&0.375e-18\end{array}\right]\alpha[m]$
$\mu^{(p,k)}[m]$ $\left[\begin{array}[]{ccc}4.4e-3&6e-3&3.5e-12\end{array}\right]\alpha_{2}[m]$
$p^{(k)}$ $\left[\begin{array}[]{ccc}2.5&2.5&0.1\end{array}\right]$
$\mu^{(k)}_{f,\ell_{1}}[m]$ $\left[\begin{array}[]{ccc}0.008&0.135&0\end{array}\right]\alpha_{2}[m]$
$\mu^{(k)}_{b,\ell_{1}}[m]$ $\left[\begin{array}[]{ccc}0&0.35&0.05\end{array}\right]\alpha_{2}[m]$
$\mu^{(k)}_{f,w-\ell_{2}}[m]$ $\frac{8e-2}{10^{-2}\times\lfloor\frac{m}{10}\rfloor+1}$
$\mu^{(k)}_{b,w-\ell_{2}}[m]$ $\frac{8e-2}{10^{-2}\times\lfloor\frac{m}{10}\rfloor+1}$
$\lambda_{E}$ $0.999999$
$\lambda_{d}$ $0.99999$
$B$ $20$
I.3.6 Hyperparameters for 3-layer CorInfoMax-EBD on MNIST and CIFAR-10 datasets for batch size 1
@@
Table 9 and 10 list the hyperparameters used in the 3-Layer CorInfoMax-EBD numerical experiments for the MNIST and CIFAR-10 datasets with a batch size of $1$, where $m$ denotes the iteration index.
Table 9: 3-Layer CorInfoMax-EBD hyperparameters: MNIST dataset, $1$-batch size.
Hyperparameter Value
$\alpha[m]$ $\frac{1}{3\times 10^{-3}\times\lfloor\frac{m}{10}\rfloor+1}$
$\alpha_{2}[m]$ $\frac{1}{3\times\lfloor\frac{m}{10}\rfloor+1}$
$\mu^{(d_{f},k)}[m]$ $\left[\begin{array}[]{ccc}4.8&3.0&5e3\end{array}\right]\alpha[m]$
$\mu^{(d_{b},k)}[m]$ $\left[\begin{array}[]{ccc}4.8&3.0&5e3\end{array}\right]\alpha[m]$
$\mu^{(d_{l},k)}[m]$ $\left[\begin{array}[]{ccc}0.0125&0.0125&0.0125\end{array}\right]\alpha[m]$ for epoch$=0$
$\left[\begin{array}[]{ccc}0.025&0.025&0.025\end{array}\right]\alpha[m]$ for epoch$>0$
$\mu^{(f,k)}[m]$ $\left[\begin{array}[]{ccc}0.11e-18&0.06e-18&0.035e-18\end{array}\right]\alpha[m]$
$\mu^{(b,k)}[m]$ $\left[\begin{array}[]{cc}1.125e-18&0.375e-18\end{array}\right]\alpha[m]$
$\mu^{(p,k)}[m]$ $\left[\begin{array}[]{ccc}2.2e-4&3e-4&3.5e-12\end{array}\right]\alpha_{2}[m]$
$p^{(k)}$ $\left[\begin{array}[]{ccc}2.5&2.5&0.1\end{array}\right]$
$\mu^{(k)}_{f,\ell_{1}}[m]$ $\left[\begin{array}[]{ccc}0.0004&0.00675&0\end{array}\right]\alpha_{2}[m]$
$\mu^{(k)}_{b,\ell_{1}}[m]$ $\left[\begin{array}[]{ccc}0&0.0175&0.0025\end{array}\right]\alpha_{2}[m]$
$\mu^{(k)}_{f,w-\ell_{2}}[m]$ $\frac{8e-2}{10^{-2}\times\lfloor\frac{m}{10}\rfloor+1}$
$\mu^{(k)}_{b,w-\ell_{2}}[m]$ $\frac{8e-2}{10^{-2}\times\lfloor\frac{m}{10}\rfloor+1}$
$\lambda_{E}$ $0.99999995$
$\lambda_{d}$ $0.99999$
$B$ $1$
Table 10: 3-Layer CorInfoMax-EBD hyperparameters: CIFAR-10 dataset, $1$-batch size.
Hyperparameter Value
$\alpha[m]$ $\frac{1}{3\times 10^{-3}\times\lfloor\frac{m}{10}\rfloor+1}$
$\alpha_{2}[m]$ $\frac{1}{3\times\lfloor\frac{m}{10}\rfloor+1}$
$\mu^{(d_{f},k)}[m]$ $\left[\begin{array}[]{ccc}4&2.5&5e3\end{array}\right]\alpha[m]$ for epoch$=0$
$\left[\begin{array}[]{ccc}16&20&5e3\end{array}\right]\alpha[m]$ for epoch$>0$.
$\mu^{(d_{b},k)}[m]$ $\left[\begin{array}[]{ccc}0&0&0\end{array}\right]\alpha[m]$
$\mu^{(d_{l},k)}[m]$ $\left[\begin{array}[]{ccc}0.025&0.025&0.025\end{array}\right]\alpha[m]$ for epoch$=0$
$\left[\begin{array}[]{ccc}0.1&0.1&0.1\end{array}\right]\alpha[m]$ for epoch$>0$
$\mu^{(f,k)}[m]$ $\left[\begin{array}[]{ccc}0.11e-18&0.06e-18&0.035e-18\end{array}\right]\alpha[m]$
$\mu^{(b,k)}[m]$ $\left[\begin{array}[]{cc}1.125e-18&0.375e-18\end{array}\right]\alpha[m]$
$\mu^{(p,k)}[m]$ $\left[\begin{array}[]{ccc}2.2e-4&3e-4&3.5e-12\end{array}\right]\alpha_{2}[m]$
$p^{(k)}$ $\left[\begin{array}[]{ccc}0.125&0.125&0.005\end{array}\right]$
$\mu^{(k)}_{f,\ell_{1}}[m]$ $\left[\begin{array}[]{ccc}0.0004&0.000675&0\end{array}\right]\alpha_{2}[m]$
$\mu^{(k)}_{b,\ell_{1}}[m]$ $\left[\begin{array}[]{ccc}0&0.0175&0.0025\end{array}\right]\alpha_{2}[m]$
$\mu^{(k)}_{f,w-\ell_{2}}[m]$ $\frac{8e-2}{10^{-2}\times\lfloor\frac{m}{10}\rfloor+1}$
$\mu^{(k)}_{b,w-\ell_{2}}[m]$ $\frac{8e-2}{10^{-2}\times\lfloor\frac{m}{10}\rfloor+1}$
$\lambda_{E}$ $0.99999995$
$\lambda_{d}$ $0.99999$
$B$ $1$
@@
I.3.7 Hyperparameters for 10-layer CorInfoMax-EBD on MNIST and CIFAR-10 datasets for batch size 1

### v3

@@ Appendix I Supplementary on numerical experiments
The models were trained on an NVIDIA Tesla V100 GPU, using the hyperparameters detailed in the sections below. Each experiment was conducted five times under identical settings, and the reported results reflect the average performance. We used the standard train/test splits for the datasets, with MNIST comprising 60,000 training examples and CIFAR-10 comprising 50,000, while both datasets included 10,000 test examples. The MNIST dataset [56] is made available under the Creative Commons Attribution-Share Alike 3.0 license. The CIFAR-10 dataset [45], originating from the University of Toronto, is publicly available for academic research purposes. Both datasets were accessed via standard deep learning library functionalities.
Rather than utilizing automatic differentiation tools, we manually implemented the gradient calculations for the EBD algorithm, utilizing batched operations to ensure computational efficiency. As a side note, the $(1-\lambda)$ factors present in the derived update expressions are absorbed into the learning rate constants and thus eliminated. In our experiments, we trained the MLP models for 120 epochs and the CNN and LC models for 100 epochs on MNIST and 200 epochs on CIFAR-10. In addition, we trained the CNN model for the CIFAR-100 dataset for 300 epochs, and the CorInfoMax-EBD (3-layer, batch size = 20) model for 60 epochs.
@@ I.1 Architectures
The architectural details of MLP, CNN and LC networks for the MNIST and CIFAR-10 datasets are shown in Tables 4 and 5, respectively, while the CNN model used in the CIFAR-100 experiments is detailed in Table 7. The structure of the MNIST and CIFAR-10 models are the same as in the reference [9], while the CIFAR-100 model closely matches [6], differing only in the MaxPool shape. In all architectures, we used ReLU as the nonlinear functions except the last layer. Furthermore, the architectural details of the biologically more realistic CorInfoMax network for MNIST and CIFAR-10 datasets are shown in Table 6. These techniques are the same as examples in Appendix J.5 of [45].
Table 4: MNIST architectures. FC: fully connected; Conv: convolutional; LC: locally connected. FC layers are reported by hidden size. Conv/LC layers are reported as (channels, kernel size, stride, padding). Pooling layers use stride 1; we report the kernel size.
@@
Table 7: CIFAR-100 architectures. Conventions are the same as in Table 4.
Convolutional
Conv1 96, $5\times 5$, 1, 2
MaxPool $2\times 2$
Conv2 128, $5\times 5$, 1, 2
MaxPool $2\times 2$
Conv3 256, $5\times 5$, 1, 2
MaxPool $2\times 2$
Dropout $p$
FC1 2048
Dropout $p$
FC2 2048
Softmax 100
I.2 CorInfoMax-EBD
In this section, we offer additional details regarding the numerical experiments conducted with the CorInfoMax Error Broadcast and Decorrelation (CorInfoMax-EBD) algorithm. Appendix I.2.1 elaborates on the general implementation details. Appendix I.2.2 presents the fundamental learning steps of the algorithm, which are based on the EBD method. Appendices I.2.3 and I.2.4 discuss the initialization of the algorithm’s variables and describe the hyperparameters. Finally, Appendix I.2.5 ( 3-Layer and batch size=20, 3-Layer batch size=1) and Appendix I.2.6 (10-Layer batch size=1) detail the specific hyperparameter configurations used in our numerical experiments for the MNIST and CIFAR-10 datasets. In Appendix I.10 we present the accuracy and loss learning curves for the CorInfoMax-EBD, shown in Figures 6.(g)-(h) and Figures 7.(g)-(h), respectively.
I.2.1 Implementation details
We implemented the CorInfoMax-EBD algorithm based on the repository available at GitHub  [^2] https://github.com/BariscanBozkurt/Supervised-CorInfoMax. This repository from Bozkurt et al. [45], used as a basis for our CorInfoMax-EBD implementation, did not specify an explicit license in its public repository at the time of access. Our use and modification are for academic research purposes, building upon the published scientific work presented in [45]. The following modifications were made to the original code:
@@ I.2.2 Algorithm
1:Learning rate parameters $\lambda_{d}$$\lambda_{E}$, $\mu^{(d,k)}[m]$, $\mu^{(f,k)}[m]$, $\mu^{(b,k)}[m]$
@@
[…]
@@
[…]
@@
[…]
@@
[…]
I.2.3 Initialization of algorithm variables
We initialize the variables ${\mathbf{W}}^{(f,k)}$, ${\mathbf{W}}^{(b,k)}$, and ${\mathbf{R}}_{{\mathbf{h}}^{(k)}\bm{\epsilon}}$ using PyTorch’s Xavier uniform initialization with its default parameters for the MNIST dataset. For the CIFAR-10 dataset is initialized with gain $0.25$. For the lateral weights ${\mathbf{B}}^{(k)}$, we first generate a random matrix ${\mathbf{J}}^{(k)}$ of the same dimensions, also using the Xavier uniform distribution, with gain=1=1 for the MNIST dataset and with gain=0.5=0.5 for the CIFAR-10 dataset. We then compute ${\mathbf{B}}^{(k)}[0]={\mathbf{J}}^{(k)}{{\mathbf{J}}^{(k)}}^{T}$, ensuring that ${\mathbf{B}}^{(k)}[0]$ is a positive definite symmetric matrix.
I.2.4 Description of hyperparameters
@@
I.2.5 Hyperparameters for 3-Layer MNIST and CIFAR-10 Models
@@
Table 9 and 10 summarizes the hyperparameters used in the 3-layer CorInfoMax-EBD experiments for the MNIST and CIFAR-10 datasets with a batch size of $20$ and $1$ respectively. The iteration index is denoted by $m$ in all expressions.
Table 9: 3-Layer CorInfoMax-EBD hyperparameters for MNIST and CIFAR-10 datasets ($B=20$).
Hyperparameter MNIST CIFAR-10
$\alpha[m]$ $\dfrac{1}{3\times 10^{-3}\times\lfloor\tfrac{m}{10}\rfloor+1}$ $\dfrac{1}{3\times 10^{-3}\times\lfloor\tfrac{m}{10}\rfloor+1}$
$\alpha_{2}[m]$ $\dfrac{1}{3\times\lfloor\tfrac{m}{10}\rfloor+1}$ $\dfrac{1}{3\times\lfloor\tfrac{m}{10}\rfloor+1}$
$\mu^{(d_{f},k)}[m]$ $96,\,60,\,1e5]\alpha[m] [$$80,\,50,\,1e5]\alpha[m]$ for epoch=0=0
$320,\,400,\,1e5]\alpha[m]$ for epoch>0>0
$\mu^{(d_{b},k)}[m]$ $96,\,60,\,1e5]\alpha[m] [$0, 0, $0,\,0]\alpha[m$]
$\mu^{(d_{l},k)}[m]$ [0.25, 0.25, $0.25,\,0.25]\alpha[m$] for epoch=0=0 [0.5, 0.5, $0.5,\,0.5]\alpha[m$] for epoch=0=0
[0.5, 0.5, $0.5,\,0.5]\alpha[m$] for epoch>0>0 [2.0, 2.0, $2.0,\,2.0]\alpha[m$] for epoch>0>0
$\mu^{(f,k)}[m]$ $\times 10^{-18},\,0.06\times 10^{-18},\,0.035\times 10^{-18}]\alpha[m] [0.11×101$$\times 10^{-18},\,0.06\times 10^{-18},\,0.035\times 10^{-18}]\alpha[m$]
$\mu^{(b,k)}[m]$ $125\times 10^{-18},\,0.375\times 10^{-18}]\alpha[m] [1.1$$125\times 10^{-18},\,0.375\times 10^{-18}]\alpha[m$]
[…]
$p^{(k)}$ [2.5, 2.5, 0.1][2.5,\,2.5,\,0.1] [2.5, 2.5, 0.1][2.5,\,2.5,\,0.1]
[…]
$\mu^{(k)}_{f,w-\ell_{2}}[m]$ $\dfrac{8\times 10^{-2}}{10^{-2}\times\lfloor\tfrac{m}{10}\rfloor+1}$ $\dfrac{8\times 10^{-2}}{10^{-2}\times\lfloor\tfrac{m}{10}\rfloor+1}$
$\mu^{(k)}_{b,w-\ell_{2}}[m]$ $\dfrac{8\times 10^{-2}}{10^{-2}\times\lfloor\tfrac{m}{10}\rfloor+1}$ $\dfrac{8\times 10^{-2}}{10^{-2}\times\lfloor\tfrac{m}{10}\rfloor+1}$
$\lambda_{E}$ $0.999999$ $0.999999$
$\lambda_{d}$ $0.99999$ $0.99999$
$m^{(d)}[m]$ $0.99\dfrac{1}{\lfloor\tfrac{m}{10}\rfloor+1}+0.999\!\left(1-\dfrac{1}{\lfloor\tfrac{m}{10}\rfloor+1}\right)$ $0.99\dfrac{1}{\lfloor\tfrac{m}{10}\rfloor+1}+0.999\!\left(1-\dfrac{1}{\lfloor\tfrac{m}{10}\rfloor+1}\right)$
$B$ 2020 2020
Table 10: 3-Layer CorInfoMax-EBD hyperparameters for MNIST and CIFAR-10 datasets ($B=1$).
Hyperparameter MNIST CIFAR-10
$\alpha[m]$ $\dfrac{1}{3\times 10^{-3}\times\lfloor\tfrac{m}{10}\rfloor+1}$ $\dfrac{1}{3\times 10^{-3}\times\lfloor\tfrac{m}{10}\rfloor+1}$
$\alpha_{2}[m]$ $\dfrac{1}{3\times\lfloor\tfrac{m}{10}\rfloor+1}$ $\dfrac{1}{3\times\lfloor\tfrac{m}{10}\rfloor+1}$
$\mu^{(d_{f},k)}[m]$ $4.8,\,3.0,\,5\times 10^{3}]\alpha[m] [$$4,\,2.5,\,5\times 10^{3}]\alpha[m]$ for epoch=0=0
$16,\,20,\,5\times 10^{3}]\alpha[m]$ for epoch>0>0
$\mu^{(d_{b},k)}[m]$ $4.8,\,3.0,\,5\times 10^{3}]\alpha[m] [$0, 0, $0,\,0]\alpha[m$]
$\mu^{(d_{l},k)}[m]$ [0.0125, 0.0125, $0.0125,\,0.0125]\alpha[m$] for epoch=0=0 [0.025, 0.025, $0.025,\,0.025]\alpha[m$] for epoch=0=0
[0.025, 0.025, $0.025,\,0.025]\alpha[m$] for epoch>0>0 [0.1, 0.1, $0.1,\,0.1]\alpha[m$] for epoch>0>0
$\mu^{(f,k)}[m]$ $\times 10^{-18},\,0.06\times 10^{-18},\,0.035\times 10^{-18}]\alpha[m] [0.11×101$$\times 10^{-18},\,0.06\times 10^{-18},\,0.035\times 10^{-18}]\alpha[m$]
$\mu^{(b,k)}[m]$ $125\times 10^{-18},\,0.375\times 10^{-18}]\alpha[m] [1.1$$125\times 10^{-18},\,0.375\times 10^{-18}]\alpha[m$]
[…]
$p^{(k)}$ [2.5, 2.5, 0.1][2.5,\,2.5,\,0.1] [0.125, 0.125, 0.005][0.125,\,0.125,\,0.005]
[…]
$\mu^{(k)}_{f,w-\ell_{2}}[m]$ $\dfrac{8\times 10^{-2}}{10^{-2}\times\lfloor\tfrac{m}{10}\rfloor+1}$ $\dfrac{8\times 10^{-2}}{10^{-2}\times\lfloor\tfrac{m}{10}\rfloor+1}$
$\mu^{(k)}_{b,w-\ell_{2}}[m]$ $\dfrac{8\times 10^{-2}}{10^{-2}\times\lfloor\tfrac{m}{10}\rfloor+1}$ $\dfrac{8\times 10^{-2}}{10^{-2}\times\lfloor\tfrac{m}{10}\rfloor+1}$
$\lambda_{E}$ $0.99999995$ $0.99999995$
$\lambda_{d}$ $0.99999$ $0.99999$
$m^{(d)}[m]$ $0.99\dfrac{1}{\lfloor\tfrac{m}{10}\rfloor+1}+0.999\!\left(1-\dfrac{1}{\lfloor\tfrac{m}{10}\rfloor+1}\right)$ $0.99\dfrac{1}{\lfloor\tfrac{m}{10}\rfloor+1}+0.999\!\left(1-\dfrac{1}{\lfloor\tfrac{m}{10}\rfloor+1}\right)$
$B$ 11 11
@@
I.2.6 Hyperparameters for 10-layer CorInfoMax-EBD on MNIST and CIFAR-10 datasets for batch size 1